Electromagnetic Waves - Physics Tutorial


PLEASE GIVE CLEAR AND ORGANIZED ANSWERS TO EACH QUESTION. IF IT ASKS YOU TO EXPLAIN, THEN CLEARLY DO SO. THANKS.


A long, thin steel wire is cut in half, and each h



A long, thin steel wire is cut in half, and each h
Answer #1

images?q=tbn:ANd9GcRtgGv487wcF4sXHfvds8m

An electromagnetic field (also EMF or EM field) is a physical field produced by electrically charged objects. It affects the behavior of charged objects in the vicinity of the field. The electromagnetic field extends indefinitely throughout space and describes the electromagnetic interaction. It is one of the four fundamental forces of nature (the others are gravitation, weak interaction and strong interaction).

The field can be viewed as the combination of an electric field and a magnetic field. The electric field is produced by stationary charges, and the magnetic field by moving charges (currents); these two are often described as the sources of the field. The way in which charges and currents interact with the electromagnetic field is described by Maxwell's equations and the Lorentz force law.

From a classical perspective in the history of electromagnetism, the electromagnetic field can be regarded as a smooth, continuous field, propagated in a wavelike manner; whereas from the perspective of quantum field theory, the field is seen as quantized, being composed of individual particles

Ot

1. Gauss's Law

\nabla \cdot \mathbf{E} = \frac {\rho}{\epsilon_0}


2. No Magnetic Monopole Law

\nabla \cdot \mathbf{B} = 0


3. Faraday's Law

\nabla \times \mathbf{E} = - \frac {\partial \mathbf{B}}{\partial t}


4. Amp